# TAPAS: Datasets for Learning the Learning with Errors Problem

### v1

@@ 4.1 Data Quality Analysis
We measure the reduction factor $\rho$ as the ratio of the means of the standard deviations of the rows of $\mathbf{R}\mathbf{A}$ and $\mathbf{A}$. A smaller $\rho$ implies that $\mathbf{RA}$ is more reduced. We preprocess each matrix $\mathbf{A}$ until $\rho$ stops decreasing. For each dataset, we report $\rho$ to 4 decimal places. In addition, per nolte2024cool , the reduction algorithms produce a cliff shape in the standard deviations of the columns of $\mathbf{R}\mathbf{A}$. Columns with standard deviations greater than $\frac{q}{\sqrt{12}}$ form the “cruel” region in the cliff because secret bits in those column indices are more difficult to recover. We also report the size of the cliff produced by the lattice reduction algorithms.
Table 3: Statistics on the datasets provided by this work. $n$ and $q$ are the LWE dimension and modulus, respectively. $\omega$ is the penalty parameter used in reduction (prior experimental work found that 10 is sufficient for Gaussian error with $\sigma=3.2$ picante ). $\rho$ is the stddev reduction, a metric reported in verde , while # cruel bits is the number of unreduced bits, used by nolte2024cool . #\# samples/matrix is slightly less than $m+n$ as we remove the rows with all zeroes. All statistics are calculated on a representative subset of 5000 examples.
$n$ $\log_{2}q$ $\omega$ $m$ $\rho$ # cruel bits # samples/matrix
$256$ 2020 $10$ 224 0.4284 37 438
$512$ $12$ $10$ 448 0.9036 411 841
$512$ $28$ $10$ 448 0.6740 225 939
$512$ $41$ $10$ 448 0.3992 69 938
$1024$ $26$ $10$ 1624 0.8600 750 $26$
@@ 4.2 Computational Cost
Here, we report the number of CPU hours required to process one matrix in each setting. The reduction time depends on the $\tau$ threshold and the size of $n$ and $m$. There is also a tradeoff between BKZ block size and CPU hours and the reduction achieved. With experimentation, we find a block size that balances both. We see that the $n=256$ dataset has the longest CPU hours per matrix, likely because it has the most aggressive reduction threshold $\tau$ relative to the problem difficulty.
Table 4: Computational cost of generating datasets provided by this work. $n$ and $q$ are the LWE dimension and modulus, respectively. $\omega$ is the penalty parameter used in reduction (prior experimental work found that 10 is sufficient for Gaussian error with $\sigma=3.2$ picante ). $\rho$ is the stddev reduction, a metric reported in verde . #\# hours/matrix is the average hours needed to reduce a matrix to the given $\tau$. Total CPU hours is hours to reduce a matrix on a single CPU times the number of matrices needed for 400M ($n=256$) or 40M ($n=512,1024$) samples.
$n$ $\log_{2}q$ $\omega$ #\# hours/matrix Total CPU Hours
$256$ 2020 10 44.31 40,486,047
$512$ $12$ 10 7.34 349,090
$512$ $28$ 10 16.88 664,971
$512$ $41$ 10 15.2 647,721
$1024$ $26$ 10 20.94 318,958
@@ 4.3 Benchmark Results
Results. Drawing on prior results wenger2024benchmarking , we generate binary secrets with varying Hamming weight and 3 or 4 cruel bits for each dataset, since these are known to be in the recoverable range for existing AI attacks. Then we run both attacks on these secrets. Out of all these experiments, we report the highest Hamming weight secret recovered for each setting. These experiments sanity check the quality of the data by showing that sparse secrets can be recovered, and they establish baseline attack performance.
Table 5: Benchmark results of ML attacks on the datasets provided by this work. $n$ and $q$ are the LWE dimension and modulus, respectively. $\omega$ is the penalty parameter used in reduction (prior experimental work found that 10 is sufficient for Gaussian error with $\sigma=3.2$ picante ). $h$ is the hamming weight of the secret (the number of non-zero elements in the secret vector of size $n$).
$n$ $\log_{2}q$ $\omega$ max $h$ recovered (binary) max $h$ recovered (ternary)
@@ 7 Future Work Leveraging Our Datasets
We would like to thank Mohamed Malhou for his valuable feedback and insights.

### v2

@@ 4.1 Data Quality Analysis
We present different statistics for each of the datasets in Table 3. We measure the reduction factor $\rho$ as the ratio of the means of the standard deviations of the rows of $\mathbf{R}\mathbf{A}$ and $\mathbf{A}$. A smaller $\rho$ implies that $\mathbf{RA}$ is more reduced. We preprocess each matrix $\mathbf{A}$ until $\rho$ stops decreasing. For each dataset, we report $\rho$ to 4 decimal places. In addition, per nolte2024cool, the reduction algorithms produce a cliff shape in the standard deviations of the columns of $\mathbf{R}\mathbf{A}$ (see Figure 2). Columns with standard deviations greater than $\frac{q}{\sqrt{12}}$ form the “cruel” region in the cliff because secret bits in those column indices are more difficult to recover. In Table 3, we also report the size of the cliff produced by the lattice reduction algorithms.
Table 3: Statistics on the datasets provided by this work. $n$ and $q$ are the LWE dimension and modulus, respectively. $m$ is the number of random indices we select at a time for subsampling. $\rho$ is the stddev reduction, a metric reported in verde, while # cruel bits is the number of unreduced bits, used by nolte2024cool. #\# samples/matrix is slightly less than $m+n$ as we remove the rows with all zeroes. All statistics are calculated on a representative subset of 5000 examples.
$n$ $\log_{2}q$ $m$ $\rho$ # cruel bits # samples/matrix
$256$ 2020 224 0.4284 37 438
$512$ $12$ 448 0.9036 411 841
$512$ $28$ 448 0.6740 225 939
$512$ $41$ 448 0.3992 69 938
$1024$ $26$ 1624 0.8600 750 $26$
@@ 4.2 Computational Cost
In Table 4, we report the number of CPU hours required to process one matrix in each setting. The reduction time depends on the $\tau$ threshold and the size of $n$ and $m$. There is also a tradeoff between BKZ block size and CPU hours and the reduction achieved. With experimentation, we find a block size that balances both. We see that the $n=256$ dataset has the longest CPU hours per matrix, likely because it has the most aggressive reduction threshold $\tau$ relative to the problem difficulty.
Table 4: Computational cost of generating datasets provided by this work. $n$ and $q$ are the LWE dimension and modulus, respectively. #\# hours/matrix is the average hours needed to reduce a matrix to the given $\tau$. Total CPU hours is hours to reduce a matrix on a single CPU times the number of matrices needed for 400M ($n=256$) or 40M ($n=512,1024$) samples.
$n$ $\log_{2}q$ #\# hours/matrix Total CPU Hours
$256$ 2020 44.31 40,486,047
$512$ $12$ 7.34 349,090
$512$ $28$ 16.88 664,971
$512$ $41$ 15.2 647,721
$1024$ $26$ 20.94 318,958
@@ 4.3 Benchmark Results
Results. Drawing on prior results wenger2024benchmarking, we generate binary secrets with varying Hamming weight and 3 or 4 cruel bits for each dataset, since these are known to be in the recoverable range for existing AI attacks. Then we run both attacks on these secrets. Out of all these experiments, we report the highest Hamming weight secret recovered for each setting in Table 5. These experiments sanity check the quality of the data by showing that sparse secrets can be recovered, and they establish baseline attack performance.
Table 5: Benchmark results of ML attacks on the datasets provided by this work. $n$ and $q$ are the LWE dimension and modulus, respectively. $h$ is the hamming weight of the secret (the number of non-zero elements in the secret vector of size $n$).
$n$ $\log_{2}q$ max $h$ recovered (binary) max $h$ recovered (ternary)
@@ 7 Future Work Leveraging Our Datasets
The authors would like to thank Mohamed Malhou for his valuable feedback and insights.